Let f:R2 to R defined by f(x)=√(x2+y2) show that f is continuous on R2

1
Expert's answer
2021-03-23T14:34:16-0400

"f(x,y)=\sqrt{(x^2+y^2)}"


now, to show that f(x,y) is continuous everywhere on "R^2" ,

we first fix an arbitrary "(a.b)" and "\u03f5>0" .


here note that "f" is indeeded the Euclidean's norm on "R^2" ,

so we can use the reverse triangle inequality.

For any "(x,y)\u2208R^2" with "|(x,y)\u2212(a,b)|<\u03b4=\u03f5"


"d((fx,y),f(a,b))=|f(x,y)\u2212f(a,b)|"

"=|\sqrt{x^2+y^2}-\sqrt{a^2+b^2}|\\
=|f(x,y)|−|f(a,b)||\\
≤|(x,y)−(a,b)|\\
<ϵ"


hence, proving "f" is continuous in our arbitrary chosen (a,b).

since, this is our arbitrary chosen ,

we conclude that "f" is continuous everywhere.

this assumes that both domain and co-domain of "f"

are equipped with corresponding euclidean metrics.


hence,proved.